# LLaMIPa: An Incremental Discourse Parser

### v1

@@ 6 Results
We first ran our finetuning experiments on the MSDC corpus, and assessed LLaMIPa’s performance on link and relation type prediction when we include the gold dialogue structure in each sample context. We then tested LLaMIPa by using the previously predicted relations in the context instead of the gold ones, as in the generation step described above, in order to have our incremental parser build a predicted structure from beginning to end. Table 2 shows that LLaMIPa3 significantly outperforms BERTLine and outperforms LLaMIPa2 both using gold structure (LLaMIPa3+g vs. LLaMIPa2+g) and predicted structure (LLaMIPa3+p vs. LLaMIPa2+p). Below we use the better performing LLaMIPa3+p for other evaluations.
To test the generalizability of the model, we also finetuned LLaMIPa3 on STAC-Sit and compared it to BERTLine on MSDC and STAC-Sit. For both link prediction and link relation prediction, LLaMIPa shows a gain of almost 10 F1 points or 11% on link and 10.5 F1 points or 15% on link relation labeling. LLaMIPa posted even more impressive gains for predictions of certain relations. When we look at F1 breakdown by relation type in Table 5, the largest relative increase comes from improved predictions for the Narration and Correction relation types, which Thompson et al., (2024) cite as the most difficult to predict given their particular characteristics.
We then compared our model to a broader range of SDRT discourse parsers on the STAC linguistic version (STAC-L), as previous work focused on this version of the corpus. Table 4 shows that LLaMIPa3 finetuned on STAC-L and evaluated on STAC-L Test scored significantly higher on link prediction and got a statistically significantly higher score on link + relation prediction over the state-of-the-art parser Chi and Rudnicky, (2022).
Nevertheless, LLaMIPa3 has substantially lower performance on STAC-L (linguistic) than on either of the situated corpora. This mirrors human efforts at annotation. Asher et al., (2016, 2021) report that annotating the STAC linguistic data set was more difficult; annotators felt that information needed to determine the discourse structure was missing.
We also investigated LLaMIPa3 for transfer learning, finetuning on STAC-L and testing on the Molweni Cleaned Test corpus of Li et al., (2024). While superior to other models, in comparison to Li et al., (2024), LLaMIPa3 fared worse on link prediction but was far superior with respect to link + relation prediction.
Overall, LLaMIPa has several advantages. It simplifies previous attach-then-type pipeline methods by predicting link and relation labels simultaneously. Previous research has given theoretical arguments that co-predicting attachment and relation labels should give better results Asher and Lascarides, (2003), as well as empirical evidence for this claim Bennis et al., (2023). Our model also outperforms prior models on all relevant corpora and pretty much all tasks we studied. We hypothesize that this performance is due to the richer contextual information supplied to LLaMIPa, a hypothesis we now explore further.
@@
| BERTLine | LLaMA3-Local | LLaMIPa3+p | LLaMIPa3+g | LLaMIPa-rand | No Structure | |
@@ 7 The role of context
A particularly remarkable demonstration of LLaMIPa’s discourse parsing prowess is its success in predicting Narration instances in the MSDC. Narrations are frequently quite long (with a significant proportion above distance 10), and so difficult to capture by local parsers like BERTLine Thompson et al., (2024). Table 3 shows Narration F1 at distances between 3 and 14 EDUs. Like other local parsers, BERTLine requires a distance cutoff that balances ignoring longer relations with mitigating the class imbalance caused by numerous unlinked candidates; for SDRT corpora, this is distance 10. As explained in Section 5, LLaMIPa allows for training up to distance 15, and impressively, it continues to make (somewhat steadily good) predictions up to distance 15.
LLaMIPa also has very strong performance in predicting instances of the Correction relation, which also has a significant number of long distance instances in the MSDC. Table 5 shows that its performance is more than double BERTLine’s on this relation; its F1 score for Correction is over 42 F1 points more than BERTLine’s.
We hypothesize that LLaMIPa has a high F1 on Narration and Correction because it has access to rich semantic information about the discourse context. It can use the meaning of previous discourse moves as well as an explicit representation of their discourse structure. This hypothesis coheres with our understanding of how Narration functions in the MSDC. Narrations connect high-level instructions between which participants clarify, question and elaborate on those instructions, as well as correct missteps (see Figure 1). While there is some statistical similarity in their surface forms, the important cues for isolating a new instruction, and hence a Narration attachment, involve certain patterns of Result and Acknowledgement links in the prior context, as well as the presence of (nonlinguistic) block placements. This pattern is regular enough that one can find some Narrations algorithmically as in Thompson et al., (2024)’s “Second Pass”: taking predictions from a first pass of the local model, Thompson et al., then apply an algorithmic approach to predict long distance Narration links from the local model’s predictions. We also tried this second pass for LLaMIPa, however there was no improvement on F1 for Narration scores, which shows that LLaMIPa had already learned this rule during training.
Besides its performance on Narration and Correction, LLaMIPa offers significant improvements over BERTLine for many other relation types, even those whose instances are typically not very long distance. Table 5 shows LLaMIPa has significantly higher scores for Result, Acknowledgment, Elaboration, Question-answer Pair, Clarification Question and Comment. We hypothesize that predictions on these too profit from complex contextual information.
LLaMIPa produces poorer results for relations such as Contrast, Alternation and Question-Elaboration. This may be because instances of these relations can be inferred reliably just from information given by the related EDUs or EEUs.
@@
First ablation: embeddings vs. context. While the previous discussion suggests that LLaMA’s larger context window is responsible for its better performance, we wanted to check this hypothesis further. So we finetuned LLaMA3-8b over pairs of EDUs as input to predict links and relation types between pairs of EDUs. This produces a local model roughly comparable to BERTLine’s, though BERTLine’s architecture is different. Table 5 shows that, compared to BERTLine, LLaMA3-Local gave stronger performance without using the prior discourse context for link and relation prediction. However, when compared to LLaMIPa3, LLaMA3-Local has significantly lower link and link+relation labeling predictions overall. LLaMIPa3’s relation predictions for many relations are also significantly better—indeed sometimes vastly superior as in the F1 for Narration. This supports our conclusion that although LLaMA’s superior embeddings play a role, many relations, and especially relations like Narration, require more contextual information like that available to LLaMIPa.
Further ablations: removing structure. Given the result of our first ablation experiment, we wanted to investigate further whether LLaMIPa was using the discourse structure for its predictions, or was solely relying on the longer input sequence of EDUs and EEUs.
@@
Our set up for LLaMIPa makes the assumption that it is the prior discourse structure that is relevant. In Table 5, there is a small but consistent drop between gold and predicted contexts in F1 performance for all relation types. This supports the conclusion that LLaMIPa takes the discourse structure of its context into account and that the quality of that structure affects prediction.
To provide more support for this conclusion, we performed another ablation in which we perturbed the structure at each increment by randomizing both the relation type and EDU index for each relation. We then tested LLaMIPa (finetuned in the usual way) on this input. LLaMIPa-rand showed a significant drop in performance from LLaMA3-Local for link and across almost all relation types (with the exception of Contrast and Question Elaboration). The drop for Narration and Correct is sharp, as well as for Question-answer Pair. We then performed a variant ablation, in which we removed all structure from LLaMIPa’s input context, leaving only a sequence of EDUs and EEUs. The performance of LLaMIPa on this data set (in the column “No Structure” of Table 5) shows another significant drop both in link and in link+relation predictions. The drop in relation labeling in addition holds for almost all relations, and a dramatic drop, especially for Narration, Correction, Clarification Question and Question-Answer Pair. This provides strong evidence that LLaMIPa needs to the use discourse structure and not just the EDU/EEU input sequence of the context to optimize its predictions for links and relation labels and that discourse structure is crucial for predicting some relations.
@@ 8 Conclusions
In this paper, we presented LLaMIPa3, an incremental SDRT style discourse parser that uses its prior discourse structure predictions to predict links and relation labels for new discourse units. We finetuned both LLaMa2-13b and LLaMA3-8b models to produce LLaMIPa2 and LLaMIPa3. We showed that LLaMIPa3 performs better than other parsers that use SDRT data sets. It excels in particular on data sets that feature conversation that contains both linguistic moves and nonlinguistic contexts that help interpret them.
We also compared LLaMIPa3 with a local model discourse parser based on LLaMA3-8b as well as on variants of LLaMIPa3 that had differing levels of access to the discourse context. LLaMIPa3 had significantly higher F1 scores for attachment and relation labeling on the data sets we used for evaluation than all of these.
The differential performance of LLaMIPa3 on STAC-L versus STAC-Sit, which reflects impressions of human annotators, suggests that nonlinguistic contexts are important for stable interpretation of situated conversation. We humans typically invent a nonlinguistic context for a conversation when one isn’t given. But such an invention is subjective and may contribute to variations in understanding and prediction. We suspect something like this might be true for LLMs too. In future work, we hope to use LLaMIPa3 to help conversational robots to improve their performance in interactive tasks.
@@
LLaMIPa, like previous parsers, presumes that the edus/eeus are already segmented. However extant work on automatic segmentation suggests that this task is relatively tractable on diverse data sets Muller et al., (2019) including even corpora of spoken conversation (Gravellier et al.,, 2021; Prevot et al.,, 2023)
Though ablations show that LLaMIPa uses discourse structure, it remains unclear how it leverages that structure to predict relations like Narration, Corrections, and QAP.
@@
Our work here has been to improve the capacities of AI models to understand the semantic structure of discourse and conversation. We see no direct ethical concerns that arise from this work, though we hope that a better grasp of the semantic structure of conversation may enable models to eventually produce less pragmatically grounded or conversational hallucinations.
@@ Appendix A Appendix: Model Parameters
Table 6 gives the hyperparameters used for finetuning LLaMIPa-2 and LLaMIPa-3 along with the computing resources. We adapted the finetuning code from the following repository444https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb. We provide LLaMIPa3 trained on MSDC as well as the MSDC test data here555https://huggingface.co/Kqte/LLaMIPa
@@ Appendix B Appendix: STAC Linguistic data
We constructed STAC-L, our STAC linguistic corpus, from the STAC data available on the STAC website (https://https://www.irit.fr/STAC/corpus.html) and adopted standard procedures for flattening CDUs. However, we found that our data set did not quite match that of Bennis et al., (2023) in terms of numbers of EDUs. Shi and Huang, (2019) also report a different number of EDUs from what we have and from what Bennis et al., (2023) report. When we finetuned and tested LLaMIPa on the STAC linguistic only data available in the BERTLine repository 666https://github.com/zineb198/LineBert, we had lower scores than what we report in Table 4. LLaMIPa-3+p had an F1 for link of 75.4 and a link+relation F1 of 57.7 on this version of STAC linguistic. On the transfer learning from STAC linguistic to Molweni Cleaned Test, LLaMIPa-3+p had an F1 of 68.5 for link and 38.7 for link+relation.

### v2

@@ 6 Results
We run finetuning experiments using the MSDC corpus, and assess Llamipa’s performance on link and link+relation prediction under the two test settings described in Section 5. We consider only the well-formed outputs during evaluation, but note that Llamipa outputs only well-formed relation instances for all experiments.
@@
Table 2 shows that Llamipa3 significantly outperforms BERTLine and outperforms Llamipa2 both using gold structure (Llamipa3+g vs. Llamipa2+g) and predicted structure (Llamipa3+p vs. Llamipa2+p).
To test the generalizability of our approach, we also finetune Llamipa on STAC-Sit and compare the Llamipa3+p outputs to BERTLine. Llamipa3+p shows a gain of F1 points on link prediction and F1 points on link+relation prediction. As mentioned in Section 5, the predicted context setting best demonstrates the desired functionality of an incremental parser. Since Llamipa3+p give the best performance in this setting, we use it for all evaluations and henceforward refer to it using the shorthand Llamipa.
Performance on STAC-L. In order to compare Llamipa to a broader range of SDRT discourse parsers, we tested it on STAC-L, as prior work has only used this linguistic version of STAC. Table 3 shows that Llamipa finetuned on STAC-L and evaluated on the STAC-L test set scores substantially higher on both link and link+relation prediction than the state-of-the-art parser Chi and Rudnicky, (2022). We note that Llamipa has a lower performance on STAC-L than on either of the situated corpora. This reflects the findings of Asher et al., (2016, 2021), who reported that annotating the STAC linguistic data set was more difficult; annotators felt that the nonlinguistic information needed to determine the discourse structure was missing (see Section 4).
| Distance | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | 11 | 12 | 13 | 14 | 15 |
We also test Llamipa on transfer learning. Following Li et al., (2024), we finetune on STAC-L and test on their Molweni-clean test set. While Llamipa outperforms all models on STAC-L, it fares worse on link prediction than Li et al., (2024) on Molweni, though it does better on link+relation prediction; a counterintuitive result given that link+relation prediction is the harder task. Taking a closer look at the link prediction results in Li et al., (2023), we see their model performs well on links of distance , but drops after that ( accuracy for distance , and for distances greater than ). Given that Molweni is a relatively simple corpus where most of the links are adjacent, their model excels on the link task. But for relation type prediction, Li et al., (2024) use the traditional method, which assigns relation types after establishing the attachment structure, and using only local information. Llamipa surpasses their model on relation type prediction, which shows that an approach predicting link and relation simultaneously performs better than a local pipeline approach, even for shallower discourse structures.
Performance on MSDC. Llamipa shows impressive gains for predictions of certain relation types. Looking at F1 breakdown by relation type in Table 5, we see the largest relative increase comes from improved predictions for Narration and Correction, which Thompson et al., (2024) cite as the most difficult to predict given that they are heavily context dependent. Narrations are frequently quite long (with a significant proportion above distance ), and difficult to capture by local parsers like BERTLine Thompson et al., (2024). Table 4 shows Narration F1 at distances between and EDUs. Like other local parsers, BERTLine requires a distance cutoff that balances ignoring longer relations with mitigating the class imbalance caused by numerous unlinked candidates; for SDRT corpora, this is distance . As indicated in Section 5, Llamipa continues to make good predictions up to distance .
Further, Llamipa is able to predict many instances of the Correction relation, which also has a number of long distance instances in the MSDC. Table 5 shows that its performance is F1 points, which is 42 points higher than and more than double that of BERTLine.
We hypothesize that Llamipa has a high F1 on Narration and Correction because it has access to rich semantic information about the discourse context: it can use the previous discourse moves as well as an explicit representation of their discourse structure. This hypothesis coheres with our understanding of how Narration functions in the MSDC. Narrations connect high-level instructions between which participants clarify, question and elaborate on those instructions, as well as correct missteps (see Figure 1). While there is some statistical similarity in their surface forms, the important cues for isolating a new instruction, and hence a Narration attachment, involve certain patterns of Result and Acknowledgement links in the prior context, as well as the presence of (nonlinguistic) block placements. This pattern is regular enough that one can find most Narrations programmatically as with Thompson et al., (2024)’s “Second Pass”. This takes predictions from a first pass of the local model for Results and Acknowledgements, then applies an algorithmic approach to predict long distance Narration links. We tried this second pass for Llamipa, however there was no improvement on F1 for Narration scores, indicating that Llamipa has already internalized this pattern during training.
Llamipa offers improvements over BERTLine for many other relation types which are typically not considered long distance. Table 5 shows Llamipa has higher scores for Result, Acknowledgment, Elaboration, Question-answer Pair, Clarification Question, and Comment. We hypothesize that predictions on these also profit from complex contextual information, something that we discuss in more detail in Section 7. Conversely, for relations where Llamipa gives slightly poorer results—Contrast, Alternation, and Question-Elaboration—we surmise that these types can be inferred reliably just from the local information given by the related EDUs or EEUs.
Overall, Llamipa has several advantages. It simplifies previous attach-then-type pipeline methods by predicting link and relation labels simultaneously. Previous research has given theoretical arguments that co-predicting attachment and relation labels should give better results Asher and Lascarides, (2003), as well as empirical evidence for this claim Bennis et al., (2023). And as we have demonstrated, Llamipa outperforms other parsers on all relevant corpora.
| BERTLine | Llama3-local | Llamipa3+p | Llamipa3+g | Llamipa+rand | Llamipa+ | |
@@ 7 The role of context
So far, the hypothesis has been that Llamipa’s state-of-the-art performance, especially for longer distance relations, is due to its ability to process richer contextual information. Whether this is a function of the superior embeddings conferred by the underlying LLM, or a function of a larger context window that allows a longer dialogue history and explicit discourse structure as prediction input, is something we explore further in this section.
@@
First ablation: embeddings vs. context size. This ablation tests whether the LLM embeddings for EDUs and EEUs account for Llamipa’s performance. We finetune Llama3-8b over pairs of elementary units as input to predict links and relation types, resulting in a local model roughly comparable to BERTLine. Table 5 shows that, compared to BERTLine, Llama3-local still performed better without using the prior discourse context for link and link+relation prediction. However, when compared to Llamipa, Llama3-local still has considerably lower predictions overall. Llamipa’s relation predictions for many relations are also better, notably for Narration. This supports our conclusion that although Llama’s superior embeddings play a role, many relations, and especially relations like Narration, require the information available in Llamipa’s larger context window.
Second ablation: perturbing input structure. Even if Llamipa requires the input available in a larger context window, it is still an open question whether it uses the discourse structure provided in this window.
@@
We already have some indication that Llamipa does use the structure when we look at Llamipa’s performance given a gold discourse structure for a context versus its predicted structure, as described in Section 5. The results of this comparison are shown in Table 5: there is a small but consistent drop between gold and predicted structure contexts in F1 performance for all relation types.
To investigate this further, we perform a second ablation where we perturb the input structure for each test sample by randomizing both the relation type and EDU index for each relation instance in the structure. When we run Llamipa on this randomized input (denoted as Llamipa+rand in Table 5) we see a drop in performance from Llamipa3+p for link and link+relation for most relation types, with a substantial drop for Narration, Correction, and Question-answer Pair.
Third ablation: removing input structure. In this ablation, we removed all structure from Llamipa’s input context, leaving only a sequence of elementary units. The performance of Llamipa on this data set (in the column Llamipa+ of Table 5) shows an even larger drop in both link and link+relation predictions compared to Llamipa+rand. Similar to Llamipa+rand, we see a substantial drop for the relation types Narration, Correction, and Question-Answer Pair. Additionally, the performance dropped drastically for Clarification Questions.
Fourth ablation: targeted perturbations. The results of the previous ablations support the hypothesis that the structure provided in the context window does have an effect on relation instance predictions. This final ablation investigates whether Llamipa predicts context dependent relations as well as it does because it leverages (or “reasons” about) the relation instances in the context structure in a manner consistent with our semantic intuitions. Here, we focus on instances of Question-answer pair (QAP) and Correction.
QAP attachments provide a straightforward example of context dependence. In the MSDC, it is common that if a prediction QAP(y,z) is made for a dialogue turn z then CLARIFQ(x,y), CONFQ(x,y), or QELAB(y,z) is in its context. That is, the questions that are answered in the QAP relations tend to be questions asked to clarify, confirm or elaborate upon the preceding context. To test whether Llamipa’s predictions conform to this pattern, we isolate the test samples where Llamipa correctly predicts QAP (Table 5), and remove all question relation types from the structure given in the context of these samples. We re-run the generation step on this amended sample set, and see a point drop in F1, from to , for QAP predictions, and no change for all other relation types (see Figure 3).
Correction relations are among the most difficult to identify in the MSDC, as discussed in Section 6. Although the language used for Corrections usually features some form of negation—No, Not like that, Oops!—-deciding where it attaches requires making inferences about how the content of the corrective move applies to a more global information state. Although this is beyond the capability of a discourse parser like Llamipa, there are still structural regularities in the MSDC that can help with retrieving Corrections.
For instance, most Corrections come about when the Builder makes the wrong moves in response to an instruction, and the Architect then gives a further instruction, which results in moves that address this by fixing the last wrong moves. In the discourse structure, these interactions create recurring patterns, “correction triangles” (Figure 4). These triangles provide a way to check Llamipa’s sensitivity to CORR(x,y) type relations in the context structure in cases where it must predict RES(y,z) CORR(x,z), the last two relations in a triangle. As with the Question-answer Pair test, we single out each test sample in which Llamipa correctly predicts the second CORR; but instead of removing CORR(x,y), we change it to ACK(x,y) (Acknowledgement(x,y)). This leaves the structure intact, but gives it a contrary sense. We re-run the generation step and see a point drop in F1, from to , for Correction predictions, and no change for all other relation types.
@@ 8 Conclusions
Llamipa is an incremental SDRT style discourse parser that uses its prior discourse structure predictions to predict links and relation labels for the discourse units in new conversation turns. As such, it can be naturally assimilated into downstream applications like conversational agents.
After finetuning both Llama2-13b and Llama3-8b models to produce Llamipa2 and Llamipa3, we show that Llamipa3 has substantially higher F1 scores than other parsers that use SDRT data sets on the link and link+relation tasks. We confirm that its performance is due to access to larger swaths of context, which are not available to parsers that use a local model. We further find evidence to support the hypothesis that Llamipa uses the explicit representations of structure in its context to predict new relation instances.
Notably, Llamipa excels on data sets that feature conversation containing both linguistic and nonlinguistic moves (actions, in the MSDC) that help interpret them. This indicates that a parser like Llamipa can furnish important discourse information to builder action models; preliminary demonstrations of this can be found in Chaturvedi et al., (2024).
@@
Llamipa, like previous parsers, presumes that the EDUs/EEUs are already segmented. However, existing work on automatic segmentation suggests that this task is relatively tractable on diverse data sets Muller et al., (2019) including even corpora of spoken conversation (Gravellier et al.,, 2021; Prevot et al.,, 2023)
@@
Our work here has been to improve the capacities of AI models to understand the semantic structure of discourse and conversation. We see no direct ethical concerns that arise from this work, though we hope that a better grasp of the semantic structure of conversation may enable the output of models to be more pragmatically grounded.
Acknowledgements
For financial support, we thank the National Interdisciplinary Artificial Intelligence Institute ANITI (Artificial and Natural Intelligence Toulouse Institute), funded by the French ‘Investing for the Future–PIA3’ program under the Grant agreement ANR-19-PI3A-000. This project has also been funded by the France 2030 program and is funded by the European Union - Next Generation EU as part of the France Relance. We also thank the projects COCOBOTS (ANR-21-FAI2-0005) and DISCUTER (ANR-21-ASIA-0005), and the COCOPIL “Graine” project of the Région Occitanie of France. This work was granted access to the HPC resources of CALMIP supercomputing center under the allocation 2016-P23060.
@@ Appendix A Appendix: Model Parameters
Table 6 gives the hyperparameters we use for finetuning Llamipa-2 and Llamipa-3 along with the computing resources. We adapt the finetuning code from the following repository333https://github.com/mlabonne/llm-course/blob/main/Fine_tune_Llama_2_in_Google_Colab.ipynb.
@@ Appendix B Appendix: STAC Linguistic data
We derive STAC-L from the STAC data available on the website (https://https://www.irit.fr/STAC/corpus.html) by adopting standard procedures for flattening CDUs. However, we find that our data set did not quite match that of Bennis et al., (2023) in terms of numbers of EDUs. Shi and Huang, (2019) also report a different number of EDUs from what we have and from what Bennis et al., (2023) report. When we finetune and test Llamipa on the STAC-L data available in the BERTLine repository 444https://github.com/zineb198/LineBert , we had lower scores than what we report in Table 3. Llamipa-3+p had an F1 for link of 75.4 and a link+relation F1 of 57.7 on this version of STAC linguistic. On the transfer learning from STAC linguistic to the Molweni-clean test set, Llamipa-3+p had an F1 of 68.5 for link and 38.7 for link+relation. We make our version of STAC-L available on the STAC corpus website cited above.